# The Boundaries of Tractability in Hierarchical Task Network Planning

### v1

@@ 4 Primitive Task Networks
We reduce from Shuffle Product with the alphabet $\Sigma={\{a,b\}}$. Construct a task for each letter in the words $u,c_{1},\ldots,c_{w}$. In particular, for each word $x\in{\{u,c_{1},c_{2},\ldots,c_{w}\}}$, let the task $t_{x,j}$ refer to its $j^{\text{th}}$ letter. Let $T_{c}$ be the set of tasks created for letters in $c_{1},\ldots,c_{w}$, and $T_{u}$ the set of tasks $t_{u,j}$, $j\in|u|$. Let $\prec^{+}$ be such that the tasks for each word form a chain in the respective order, i.e., let $\prec$ consist of the edges $(t_{x,j},t_{x,j+1})$ for each $x\in{\{u,c_{1},\ldots,c_{w}\}}$ and $j\in[|x|-1]$. Further, add a task $t_{g}$ and let $(t_{u,|u|},t_{g})\in\prec$. Then, $\prec^{+}$ partitions $T$ into $w+1$ chains (see Fig. 4).
@@
Table 1: Preconditions and effects of actions in the proof of Thm. 4.
@@
Figure 4: Constructed task network for reductions from Shuffle Product with words $c_{1},\ldots,c_{w}$ and $u$. Each word is represented by a chain of tasks where each task represents a letter in that word. The colors orange and blue represent the two different letters in the alphabet. For the reduction to Plan Validation (Theorem 3), the task network only consists of vertices on the LEFT side. For the reductions to State Reachability and Plan Existence (Theorem 4), a task’s action is defined by its color and whether it is on the LEFT or RIGHT side.
Having established the intractability of all three problems on collections of paths, it is natural to wonder whether State Reachability and Plan Existence can at least be solved on collections of stars. Here, too, we provide a negative answer via a reduction from the strongly NP-hard 3-Partition problem Garey and Johnson (1979):
3-Partition
Input: $m,B\in\mathbb{N}$ and $3m$ positive integers $n_{1},\ldots,n_{3m}$ such that $\sum_{i=1}^{3m}n_{i}=mB$.
Question: Can the integers $n_{i}$ be partitioned into $m$ triples such that each triple sums up to $B$?
Theorem 5.
State Reachability and Plan Existence are NP-hard on primitive task networks that are forests of stars, even if $|A|=8$ and the state transition graph has at most $9$ states.
Proof.
We reduce from 3-Partition. We employ the set of actions $A={\{a_{\emptyset},a_{\beta},a_{\hat{\beta}},a_{\gamma},a_{\hat{\gamma}},a_{\delta},a_{\hat{\delta}},a_{g}\}}$. The preconditions and effects of these actions are described in Figure 5, which also illustrates the construction of the following task network. For each $i\in[3m]$, create a star representing the integer $n_{i}$ as follows. Create a task $t_{i}$ and two task sets $T_{i},T_{i}^{\prime}$, each of cardinality $n_{i}$. Let $\alpha(t_{i})=a_{\emptyset}$. For all $t\in T_{i}$, let $\alpha(t)=a_{\hat{\beta}}$ and $(t,t_{i})\in\prec$. For all $t\in T_{i}^{\prime}$, let $\alpha(t)=a_{\gamma}$ and $(t_{i},t)\in\prec$. Next, create $m$ more gadgets representing the $m$ triples to be built, where each triple is represented by a star as follows. For all $i\in[m-1]$, the $i^{\text{th}}$ triple has a center vertex $t_{\delta,i}$, a set $T_{\hat{\gamma}}$ of $B$ tasks, and a set $T_{\beta}$ of $B$ tasks. Let $\alpha(t_{\delta,i})=a_{\delta}$. For all $t\in T_{\hat{\gamma}}$, let $\alpha(t)=a_{\hat{\gamma}}$ and $(t,t_{\delta,i})\in\prec$. For all $t\in T_{\beta}$, let $\alpha(t)=a_{\beta}$ and $(t_{\delta,i},t)\in\prec$. The star representing the last triple differs from the above construction. The last triple has a task $t_{\delta,m}$ and a set $T_{\hat{\gamma}}$ of $B$ tasks as described above, but no set $T_{\beta}$. To initiate things, there is also an independent set of $B$ tasks all with action $a_{\beta}$. Lastly, create a star to count the completed triples. This star consists of a center vertex $t_{g}$ and a set $T_{\hat{\delta}}$ of $m$ leaf-tasks. Let $\alpha(t_{g})=a_{g}$ and, for each $t\in T_{\hat{\delta}}$, let $\alpha(t)=a_{\hat{\delta}}$ and $(t,t_{g})\in\prec$. Let the initial state be $s_{0}={\{\hat{\beta},\hat{\gamma}\}}$. For State Reachability, let $s_{g}={\{\text{GOAL}\}}$. We prove that these instances are yes-instances if and only if the 3-Partition instance is a yes-instance.
$a_{\emptyset}$emptyset $\emptyset$ $\emptyset$
$a_{\beta}$${\{\hat{\beta}\}}$${\{\hat{\beta}\}}$${\{\beta\}}$
$a_{\hat{\beta}}$${\{\beta\}}$${\{\beta\}}$${\{\hat{\beta}\}}$
$a_{\gamma}$${\{\hat{\gamma}\}}$${\{\hat{\gamma}\}}$${\{\gamma\}}$
$a_{\hat{\gamma}}$${\{\gamma\}}$${\{\gamma\}}$${\{\hat{\gamma}\}}$
$a_{\delta}$emptyset $\emptyset$ ${\{\delta\}}$
$a_{\hat{\delta}}$${\{\delta\}}$${\{\delta\}}$emptyset
[…]
Figure 5: Actions and constructed task network for the reduction in the proof of Theorem 5. The label at each task $t$ either refers to both its name and action ($t:\alpha(t)$) or, for legibility, just its action.
Suppose there is a 3-Partition of the integers $n_{i}$. We build a task sequence that enumerates these triples. Let $(n_{x},n_{y},n_{z})$ be the first triple in the solution. Initially, there are $B$ tasks with action $a_{\beta}$ that are executable. Alternate executing these tasks and leaf-tasks with action $a_{\hat{\beta}}$ of the stars with centers $t_{x}$, $t_{y}$, and $t_{z}$. This is possible as $n_{x}+n_{y}+n_{z}=B$ and the preconditions and effects of $a_{\beta}$ and $a_{\hat{\beta}}$ enforce this alternating behavior. Now, execute $t_{x}$, $t_{y}$, and $t_{z}$. Then, alternate between executing one of the $B$ tasks with action $a_{\gamma}$ in the stars of $t_{x}$, $t_{y}$, and $t_{z}$, and the $B$ tasks with action $a_{\hat{\gamma}}$ in the star of $t_{\delta,1}$. Then, execute $t_{\delta,1}$, followed by one of the not yet executed tasks of action $t_{\hat{\delta}}$. After executing this sequence, we are in a similar state as in the beginning. Once more, there are $B$ tasks with action $a_{\beta}$ executable. However, one of the tasks with action $a_{\hat{\delta}}$ and all of the tasks in the star with center $t_{\delta,1}$, as well as the three stars corresponding to $n_{x}$, $n_{y}$, and $n_{z}$ are executed. After repeating the above procedure for all $m$ triples, the only task remaining is $t_{g}$ (since the last star representing a triple does not have a set $T_{\beta}$) and it is executable, which results in a yes-instance for State Reachability and Plan Existence.
For the other direction, suppose we have a yes-instance of State Reachability or Plan Existence. This implies that there is a sequence executing $t_{g}$. This task is only executable after the execution of all $m$ tasks with action $a_{\hat{\delta}}$, which in turn requires the execution of all $m$ tasks $t_{\delta,i}$ for $i\in[m]$. Without loss of generality, let these tasks be executed in the order $t_{\delta,1},\ldots,t_{\delta,m}$. We define multisets $S_{i}$ for $i\in[m]$ as follows. Let $S_{1}={\{n_{j}\mid t_{j}\text{ was executed before }t_{\delta,1}\}}$ and, for $i\in[m-1]$, let $S_{i+1}={\{n_{j}\mid t_{j}\text{ was executed after $t_{\delta,i}$ and before }t_{\delta,i+1}\}}$. Observe that $P={\{S_{i}\mid i\in[m]\}}$ partitions the integers $n_{1},\ldots,n_{3m}$. In order to execute a task $t_{j}$ for $j\in[3m]$, first its $n_{j}$ in-neighbors and, by their preconditions, $n_{j}$ tasks of action $a_{\beta}$ have to be executed. The execution of the task $t_{j}$ then enables the execution of $n_{j}$ tasks of action $a_{\gamma}$. Note that, for the execution of $t_{\delta,1}$, exactly $B$ tasks with action $a_{\beta}$ are available, and hence, $\sum_{n\in S_{1}}\leq B$, and the execution of $B$ tasks with action $a_{\gamma}$ is required, and hence, $\sum_{n\in S_{1}}\geq B$. Thus, $\sum_{n\in S_{1}}=B$ and, after the execution of $t_{\delta,1}$, there are once more exactly $B$ tasks of action $a_{\beta}$ available and each of the not yet executed tasks $t_{j}$ still has $n_{j}$ not yet executed in-neighbors. Hence, we can repeatedly apply our argument and obtain $\sum_{n\in S_{i}}=B$ for all $i\in[m]$. Thereby, $P$ is a 3-Partition. Lastly, observe that, for each reachable state $s$ for which $\text{GOAL}\notin s$, either $\beta\in s$ or $\hat{\beta}\in s$, either $\gamma\in s$ or $\hat{\gamma}\in s$, and either $\delta\in s$ or $\delta\notin s$. Since $s={\{\text{GOAL}\}}$ when $\text{GOAL}\in s$, this gives $k\leq 2^{3}+1=9$ reachable states. ∎
Chains and Antichains.
Given the seemingly discouraging lower bounds obtained so far in this section, one might rightfully wonder whether there are reasonable structural restrictions on the network which can be exploited to obtain polynomial-time algorithms. In this subsection, we chart the route towards identifying these by first establishing the tractability of all considered problems (under the state-space restrictions justified in Section 3) in the extremes of $\prec^{+}$, that is, where $T$ is a chain (i.e., total order) or an antichain ($\prec^{+}\ =\emptyset$).
We begin by showing that the problems are easily solved on chains. Recall from our discussion in Section 2 that on primitive task networks, polynomial-time solvability results for Action Executability carry over to Plan Existence.
Observation 6.
Plan Validation, State Reachability, and Action Executability are linear-time solvable on primitive task networks such that $\prec^{+}$ defines a total order.
Proof.
Note that $\prec^{+}$ defines the exact order in which the tasks in $T$ must be executed and that in order to execute a certain task, all tasks that precede it in the order must be executed first. It thus suffices to iteratively execute the tasks in the order given by $\prec^{+}$ and at each step check whether the respective action given by $\alpha$ can be executed in the current state and otherwise abort. For State Reachability and Action Executability, we have a yes-instance if after some step the target state is reached or the action set is covered. For Plan Validation with plan $p$, we have a yes-instance if, for the $i^{\text{th}}$ task $t_{i}$, we have $\alpha(t_{i})=p[i]$ for all $i\in[|T|]$. ∎
It is also trivial to solve Plan Validation on antichains.
Observation 7.
Plan Validation is linear-time solvable on primitive task networks such that $\prec^{+}=\emptyset$.
Proof.
Since $\prec^{+}=\prec=\emptyset$, there are no precedence constraints on the order the tasks may be executed with respect to task names, and thus, we can simply map the tasks to their corresponding multiset of actions. Deciding whether a plan is executable in the task network is achieved by first checking whether the number of action occurrences in the sequence matches the multiset of actions and then simulating the sequence to check whether the preconditions of each action are fulfilled. ∎
Handling antichains for the remaining problems already requires more effort. For State Reachability, we can obtain an algorithm via an analysis of how a hypothetical solution traverses the state graph.
Theorem 8.
State Reachability is polynomial-time solvable on primitive antichain task networks with constantly many states.
Proof.
Let $k$ be the number of states, $s_{0}$ be the initial state and $s_{g}$ the target state. We begin by computing the state transition graph ${\cal G}$ of ${\cal I}$ in at most quadratic time. In ${\cal G}$, if there are more than $k$ edges between any pair of states, arbitrarily select $k$ of them and discard the rest. Then, consider all simple edge paths from $s_{0}$ to $s_{g}$. Note that there are $\mathcal{O}\!\left(2^{k}\cdot k^{k-1}\right)$ simple edge paths, as there are $\mathcal{O}\!\left(2^{k}\right)$ simple vertex paths, each of length at most $k-1$, and each step might choose between up to $k$ different edges. For each simple edge path, check whether there is any action that is used more times in the path than there are tasks of that action in $T$. If there exists a simple edge path for which the number of required tasks does not exceed the number of available ones, then give a positive answer, and otherwise return that it is a no-instance.
If such a path is found, the instance is indeed a yes-instance as there is a task sequence that results in a path to $s_{g}$ in ${\cal G}$. For the other direction, let there be a task sequence that leads to the state $s_{g}$. Consider the subsequence obtained by exhaustively removing cycles in ${\cal G}$ from that sequence. The resulting task sequence describes a simple edge path from $s_{0}$ to $s_{g}$. It might employ edges that were discarded from ${\cal G}$ by the algorithm. However, in this case there are $k$ other edges connecting the respective states in ${\cal G}$. Thus, one can choose to replace deleted edges by existing ones in such a way that each replacement action occurs at most once in the path, which ensures that there are enough tasks to actually walk this path. Hence, there is a path from $s_{0}$ to $s_{g}$ that only uses edges that are present in ${\cal G}$ after the deletion and for which the number of required tasks does not exceed the number of available ones. As the algorithm will find this path, it will correctly give a positive output.
Reducing the number of edges in ${\cal G}$ takes $\mathcal{O}\!\left(|A|\cdot k\right)$ time, while the number of tasks for each action can be counted in $\mathcal{O}\!\left(|T|\right)$ time. For each path of length at most $k$, it takes $\mathcal{O}\!\left(k\right)$ time to check whether there are enough tasks for each action in the path. This gives a total runtime of $\mathcal{O}\!\left(|{\cal I}|^{2}+|A|k+|T|+2^{k}\cdot k^{k}\right)$, where the middle two terms are always dominated by either the first or the last term. Hence, any instance ${\cal I}$ of State Reachability in which the task network is primitive and $\prec^{+}=\emptyset$ can be solved in $\mathcal{O}\!\left(|{\cal I}|^{2}+2^{k}\cdot k^{k}\right)$ time, where $k$ is the number of states. ∎
Observe that when trying to adapt the above approach for Plan Existence or, more generally, for Action Executability, we might now have to visit a state multiple times as multiple tasks or actions might be executable in that state alone. The number of times a state is visited is bounded by the size of the action set $S$, however this insight only yields a polynomial-time algorithm if $k+|S|$ is fixed. In particular, we note that for the reduction from Plan Existence to Action Executability, it may possibly happen that $|S|=\Theta(|T|)$. Nevertheless, we obtain a polynomial-time algorithm for both problems on instances with fixed $k$ by a more complex approach which exploits the fact that sufficiently long walks in the state transition graph will revisit the same simple cycle multiple times.
Before we present the algorithms, we will need to introduce two reduction rules and the notion of equivalent actions.
Definition 4.1.
For an HTN planning instance with state transition graph ${\cal G}=({\cal S},{\cal T},\beta)$, let $a_{1},a_{2}\in A$ and, for all $s,s^{\prime}\in{\cal S}$, let $(a_{1},s,s^{\prime})\in{\cal T}$ if and only if $(a_{2},s,s^{\prime})\in{\cal T}$. Then, $a_{1}$ and $a_{2}$ are equivalent actions.
@@
We note that there are at most $(k+1)^{k}$ equivalence classes because each class is identified by a vector $D$ of length $k=|{\cal S}|$ over ${\cal S}\cup{\{\emptyset\}}$. Thereby, $D[i]=s$ is to be interpreted as the equivalence class transforms the $i^{\text{th}}$ state into state $s$, and $D[i]=\emptyset$ implies the class does not transform the $i^{\text{th}}$ state into anything as its preconditions are not fulfilled. This allows us to define a reduction rule reducing the number of different actions. A reduction rule is safe if the rule preserves (or correctly identifies) yes- and no-instances.
@@
For any primitive task network $tn=(T,\prec^{+},\alpha)$ and action $a\in A$, let $T_{a}={\{t\in T\mid\alpha(t)=a\}}$ and $\max(a)=|T_{a}|$. For Action Executability instances with action multiset $S$, let $\min(a)$ denote the number of occurrences of $a$ in $S$. Note that a primitive Action Executability instance has a solution if and only if there is a task sequence that is executable in $tn$ and $s_{0}$ in which each action $a$ occurs at least $\min(a)$ and at most $\max(a)$ times. We now provide the reduction rules we will use in our algorithm for Action Executability.
@@
Consider a primitive Action Executability instance such that Reduction Rule R0 cannot be applied. Let $a_{1},a_{2}\in A$ with $a_{1}\neq a_{2}$ be equivalent actions. If, for all $t_{1}\in T_{a_{1}},t_{2}\in T_{a_{2}},t\in T$, the following holds:

### v2

@@ 4 Hierarchical Task Networks
pairwise non-isomorphic task networks. ∎
Having defined our measures of interest, we proceed to formalize the set of problems our meta-theorem will capture. Intuitively, the aim is to formalize the set of all problems which deal with compound HTNs by decomposing them.
Definition 2.
Let PR be an arbitrary HTN problem. Then, PR is decomposable if and only if, for all non-primitive initial task networks $tn$, PR on $tn$ is a yes-instance if and only if there is a primitive task network $tn^{\prime}$ such that $tn$ can be decomposed into $tn^{\prime}$ and $tn^{\prime}$ is a yes-instance of PR.
Note that all problems discussed in this work are decomposable. Next, in order to capture the general notions of “tractability” for various HTN problems on networks where certain measures are bounded, we need to slightly restrict the measures in question to avoid entirely degenerate measures.
Definition 3.
Let $\kappa$ be any numerical measure of HTNs in a domain $D$, i.e., a mapping that assigns each HTN a non-negative number. We say that $\kappa$ is stable if there exists a computable function $f$ with the following property: for each primitive HTN $tn^{*}$ consisting of $c$ tasks and each HTN $tn$ containing some compound task $t$, the HTN $tn^{\prime}$ obtained from $tn$ by decomposing $t$ into $tn^{*}$ satisfies $\kappa(tn^{\prime},D)\leq f(\kappa(tn,D),c)$.
We say that a stable measure $\kappa$ is a stable tractability measure for a decomposable HTN problem PR if, for each constant $c$, PR can be solved in polynomial time on all primitive HTNs $tn$ such that $\kappa(tn,D)\leq c$. Notably, the (generalized) partial order width measure fits in that definition, where its stability is due to the fact that the addition of $c$ vertices to a graph can only increase its (generalized) partial order width by at most $c$.
Observation 1.
(Generalized) partial order width is a stable tractability measure for Plan Verification. The sum of the (generalized) partial order width and the number of states in the state space is a stable tractability measure for Action Executability and State Reachability.
We are now ready to establish the first meta-theorem.
Theorem 8.
Let PR be a decomposable HTN problem and $\kappa$ a stable tractability measure. Then, PR is polynomial-time solvable when restricted to any initial network $tn$ in domain $D$ where ${C_{d}},{C_{\#}},{C_{s}},$ and $\kappa(tn,D)$ are constants.
Proof.
Let ${\cal I}$ have task network $tn=(T,\prec^{+},\alpha)$. If there are no compound tasks in $T$, then we apply an algorithm for primitive networks. Thus, assume ${C_{d}}\geq 1$. Then, PR in $tn$ can be solved by listing all the primitive instances $tn$ can be decomposed into and solving PR in each of these. By Lemma 7, there are at most ${C_{c}}^{g({C_{d}},{C_{\#}},{C_{s}})}\leq|{\cal I}|^{g({C_{d}},{C_{\#}},{C_{s}})}$ such instances ${\cal I}_{p}$ for $g({C_{d}},{C_{\#}},{C_{s}})=\sum_{i=0}^{{C_{d}}-1}{C_{\#}}\cdot{C_{s}}^{i}$ and, for each of these, $|{\cal I}_{p}|\leq|{\cal I}|+f^{\prime\prime}({C_{d}},{C_{\#}},{C_{s}})$ for some computable function $f^{\prime\prime}$. Furthermore, for each initial task network $tn_{p}$ of an instance ${\cal I}_{p}$, we have $\kappa(tn_{p},D)\leq f^{\prime\prime\prime}(\kappa(tn,D),{C_{d}},{C_{\#}},{C_{s}})$, for some computable function $f^{\prime\prime\prime}$ as $\kappa$ is a stable measure and $tn_{p}$ is obtained from $tn$ by sequentially decomposing ${C_{\#}}$ compound tasks into a total of $f^{\prime\prime}({C_{d}},{C_{\#}},{C_{s}})$ primitive tasks. Thus, if primitive instances ${\cal I}_{p}$ can be solved in time at most $|{\cal I}_{p}|^{f(\kappa(tn_{p},D))}$, then the total runtime (for some computable function $f^{\prime}$) is
$\begin{split}\mathcal{O}\Big{(}(|{\cal I}|+f^{\prime\prime}({C_{d}},{C_{\#}},{C_{s}}))^{f(f^{\prime\prime\prime}(\kappa(tn,D),{C_{d}},{C_{\#}},{C_{s}}))}\\ \cdot{C_{c}}^{g({C_{d}},{C_{\#}},{C_{s}})}\Big{)}={\mathcal{O}\!\left(|{\cal I}|^{f^{\prime}(\kappa(tn,D),{C_{d}},{C_{\#}},{C_{s}})}\right)}.\qed\end{split}$
We complement Theorem 8 via lower bounds which show that bounding $C_{d}$, $C_{\#}$, and $C_{s}$ is necessary. Specifically, we prove that if any one of these measures is not bounded by a constant, then some of the considered problems become NP-hard on compound networks, even for settings that were easily solvable on primitive networks. While other reductions show the hardness deriving from decomposing non-primitive networks (see, e.g., Erol et al. (1996); Alford et al. (2015a); Behnke et al. (2015)), our reduction is stronger in the sense that it bounds all but any one of the parameters $C_{d}$, $C_{\#}$, $C_{s}$.
Theorem 9.
For each $\iota\in\{{C_{d}},{C_{\#}},{C_{s}}\}$, Plan Existence, Action Executability, and State Reachability remain NP-hard when restricted to instances which have (generalized) partial order width at most $1$ and where $\{{C_{d}},{C_{\#}},{C_{s}},{C_{c}}\}\setminus\{\iota\}$ are upper-bounded by a constant.
Proof Sketch.
We reduce from Multicolored-Clique, which is NP-hard and asks whether a given properly $k$-vertex-colored graph $G$ contains a $k$-clique Fellows et al. (2009). The reduction creates a compound task for each vertex and edge of $G$. These can all either be decomposed into a primitive vertex- or edge-task, or an empty “dummy” network. The actions and state space of the output instance are set up in a way which ensures that, for each color class, only a single primitive vertex task can be executed. The vertices for which we execute the corresponding primitive vertex task are considered as “chosen”, and a primitive edge task can only be executed if both of its incident vertices are chosen. Based on the targeted HTN problem, we ensure the instance admits a solution if and only if it is possible to execute ($k\choose 2$ primitive edge tasks, matching each of the ($k\choose 2$ pairs of colors.
The above completes the reduction for the case where $\iota={C_{\#}}$. For the case where $\iota={C_{s}}$, we slightly adapt the construction by having the whole network (as described above) be the result of decomposing a single initial compound task $t_{0}$. Finally, when $\iota={C_{d}}$ we build on the case of $\iota={C_{s}}$ by making the decomposition from $t_{0}$ “gradual” as opposed to only occurring in a single step. ∎
5 A Parameterized Analysis of HTNs
In this final section, we ask whether—and for which parameterizations—our problems on primitive HTNs admit fixed-parameter algorithms. While there are many graph-theoretic parameters typically used to achieve fixed-parameter tractability, none of the “usual suspects” in that regard will help with our problems due to the hardness on forests of stars Lin and Bercher (2023) and sets of chains (Theorems 1 and 2). Indeed, this applies to treewidth Robertson and Seymour (1986), the feedback edge number Ganian and Korchemna (2021), and similar measures defined on digraphs Ganian et al. (2016).
On the other hand, our problems are polynomial-time solvable for constant generalized partial order width and so we need a more refined notion of hardness to rule out fixed-parameter tractability in this case. In particular, hardness w.r.t. a parameter $k$ for the complexity classes ${\textsf{W}}[1]$ or ${\textsf{W}}[2]$ rules out any fixed-parameter algorithm under the well-established assumption that ${\textsf{W}}[1]\neq{\textsf{FPT}}$. As the Shuffle Product problem is W[2]-hard when parameterized by the width $w$ van Bevern et al. (2016), Theorems 1 and 2 also imply the W[2]-hardness of the considered problems.
Fixed-Parameter Tractability via Vertex Cover.
We contrast these lower bounds with a fixed-parameter algorithm for primitive networks that exploits a different parameterization, namely the vertex cover number (vcn) of $G_{\prec}$. A vertex cover of $G_{\prec}$ is a subset of tasks $V\subseteq T$ such that, for all $t\prec t^{\prime}$, $t\in V$ and/or $t^{\prime}\in V$; the vcn of $G_{\prec}$ is the minimum size of a vertex cover of $G_{\prec}$. While the vcn has been used to obtain fixed-parameter algorithms for many other problems Balko et al. (2022); Blazej et al. (2023); Bodlaender et al. (2023); Chalopin et al. (2024), the techniques from previous works do not easily translate to HTN problems. Instead, our algorithms use a delicate branching routine whose correctness requires a careful analysis of the state space.
Theorem 10.
Plan Verification is fixed-parameter tractable on primitive task networks parameterized by the vcn of $G_{\prec}$.
Proof Sketch.
We must order the tasks so that they respect $\prec$ and correspond to the input plan $p$. We branch over all orderings of the tasks in a minimal vertex cover $V$ and discard those which do not respect $\prec$. Let the priority of a task be higher the earlier one of its out-neighbors in $V$ is placed in the ordering. Tasks in $V$ have priority just below the last of their in-neighbors, and tasks outside $V$ with no out-edges have no priority. We iterate through $p$ and, at each step, select a task for that position that has the respective action and has all of its predecessors in $\prec$ already selected. If there are multiple candidates, we choose a task with highest priority, ensuring that order constraints are met as early as possible. ∎
Theorem 11.
State Reachability and Action Executability on primitive networks are fixed-parameter tractable when parameterized by the vcn of $G_{\prec}$ plus the number of states in the state transition graph.
Proof Sketch.
We branch over the order of vertices in the vertex cover as in Theorem 10, but need more involved arguments for constructing the task sequence. Inspired by the proof of Theorem 4, we carefully branch on which paths and cycles the sequence takes in the state transition graph in between visiting the tasks of the vertex cover, and use an ILP to determine the number of times each cycle is traversed. ∎
A Meta-Theorem for Fixed-Parameter Tractability.
As our final contribution, we show that the algorithmic meta-theorem for polynomial-time solvability can be lifted to the setting of fixed-parameter tractability if we also restrict the maximum “breadth” of a compound task, i.e., ${C_{c}}$. To do this, we first need to define a suitable notion of stability. We call $\kappa$ a stable fixed-parameter tractability measure for a decomposable HTN problem PR if $\kappa$ is a stable measure and PR is fixed-parameter tractable on primitive task networks parameterized by $\kappa$. The vcn is an example of this notion as the addition of $c$ vertices to a graph increases its vcn by at most $c$.
Observation 2.
The vcn is a stable fixed-parameter tractability measure for Plan Verification. The sum of the vcn and the number of states in the state space is a stable fixed-parameter tractability measure for Action Executability and State Reachability.
We now establish the parameterized analog of Theorem 8, which extends Theorems 10 and 11 to compound networks:
Theorem 12.
Let PR be a decomposable HTN problem and $\kappa$ a stable fixed-parameter tractability measure. If ${C_{d}}\neq\infty$, then PR is fixed-parameter tractable by the combined parameters ${C_{c}},{C_{d}},{C_{\#}},{C_{s}},$ and $\kappa(tn,D)$ for the initial network $tn$.
Proof.
Let $tn=(T,\prec^{+},\alpha)$. If there are no compound tasks in $T$, i.e., ${C_{d}}=0$, the statement immediately holds by applying an algorithm for primitive networks. Thus, assume ${C_{d}}\geq 1$. Then, PR in $tn$ is solved by listing all primitive instances $tn$ can be decomposed into and solving PR in each of these. By Lemma 7, there are at most ${C_{c}}^{\sum_{i=0}^{{C_{d}}-1}{C_{\#}}\cdot{C_{s}}^{i}}$ such instances ${\cal I}_{p}$ and, for each of these, $|{\cal I}_{p}|\leq|{\cal I}|+f^{\prime\prime}({C_{d}},{C_{\#}},{C_{s}})$ for some computable function $f^{\prime\prime}$. Further, for each initial task network $tn_{p}$ of an instance ${\cal I}_{p}$, we have $\kappa(tn_{p},D)\leq f^{\prime\prime\prime}(\kappa(tn,D),{C_{d}},{C_{\#}},{C_{s}})$, for some computable function $f^{\prime\prime\prime}$ as $\kappa$ is a stable measure. Thus, if primitive instances ${\cal I}_{p}$ can be solved in time $f(\kappa(tn_{p},D))|{\cal I}_{p}|^{\mathcal{O}\!\left(1\right)}$, then the total running time can be upper-bounded by ${\mathcal{O}\!\left({C_{c}}^{g({C_{d}},{C_{\#}},{C_{s}})}\cdot f^{\prime}(\kappa(tn,D),{C_{d}},{C_{\#}},{C_{s}})\cdot|{\cal I}|^{{\mathcal{O}\!\left(1\right)}}\right)}$ for computable functions $g$ and $f^{\prime}$. ∎
6 Concluding Remarks
This article provides a comprehensive understanding of the complexity-theoretic landscape of several fundamental problems on HTNs. Our results include strong algorithmic lower bounds and complementary positive results—not only for specific problems, but also in the form of meta-theorems that can be used for other HTN problems. For further work, while we have provided lower bounds justifying all the restrictions applied in the algorithmic meta-theorem for polynomial-time solvability (Theorem 8), in the more refined fixed-parameter tractability setting we leave it open whether the “hierarchical breadth” ${C_{c}}$ needs to be part of the parameterization in order to establish Theorem 12.
6.0.1 Acknowledgements
This research was funded in whole or in part by the Austrian Science Fund (FWF) [10.55776/Y1329].
Appendix A Appendix: Proof Details
Proof of Theorem 2.
@@ Appendix A Appendix: Proof Details
We reduce from Shuffle Product with the alphabet $\Sigma={\{a,b\}}$. Construct a task for each letter in the words $u,c_{1},\ldots,c_{w}$, i.e., for each word $x\in{\{u,c_{1},c_{2},\ldots,c_{w}\}}$, let the task $t_{x,j}$ refer to its $j^{\text{th}}$ letter. Let $T_{c}$ be the set of tasks created for letters in $c_{1},\ldots,c_{w}$, and $T_{u}$ the set of tasks $t_{u,j}$, $j\in|u|$. Let $\prec^{+}$ be such that the tasks for each word form a chain in the respective order, i.e., let $\prec$ consist of the edges $(t_{x,j},t_{x,j+1})$ for each $x\in{\{u,c_{1},\ldots,c_{w}\}}$ and $j\in[|x|-1]$. Further, add a task $t_{g}$ and let $(t_{u,|u|},t_{g})\in\prec$. Then, $\prec^{+}$ partitions $T$ into $w+1$ chains (see Fig. 3).
@@
Table 1: The actions in the proof of Thm. 2.
@@
Action Executability on primitive Antichains (Thm. 4).
@@
Let $k=O(1)$ be the number of states in the state transition graph. We say that two actions $a_{1},a_{2}$ are equivalent if they are executable from the same states and lead to the same state, i.e., for all states $s,s^{\prime}$, there is an edge (with label) $a_{1}$ from $s$ to $s^{\prime}$ if and only if there is an edge $a_{2}$ from $s$ to $s^{\prime}$ in the state transition graph. We note that there are at most $(k+1)^{k}$ equivalence classes because each class is identified by a vector $D$ of length $k=|{\cal S}|$ over ${\cal S}\cup{\{\emptyset\}}$. Thereby, $D[i]=s$ is to be interpreted as the equivalence class transforms the $i^{\text{th}}$ state into state $s$, and $D[i]=\emptyset$ implies the class does not transform the $i^{\text{th}}$ state into anything as its preconditions are not fulfilled. This allows us to define a reduction rule reducing the number of different actions. Reduction rules are safe if they preserve yes- and no-instances.
@@
For a primitive task network $tn=(T,\prec^{+},\alpha)$ and action $a\in A$, let $T_{a}={\{t\in T\mid\alpha(t)=a\}}$ and $\max(a)=|T_{a}|$. For Action Executability instances with action multiset $S$, let $\min(a)$ be the number of occurrences of $a$ in $S$. A primitive Action Executability instance is a yes-instance if and only if there is a task sequence that respects $\prec^{+}$, is executable in $s_{0}$ and in which each action $a$ occurs at least $\min(a)$ and at most $\max(a)$ times. We now provide two reduction rules for Action Executability.
@@
Consider a primitive Action Executability instance where Reduction Rule R0 cannot be applied. Let $a_{1},a_{2}\in A$ with $a_{1}\neq a_{2}$ be equivalent actions. If, for all $t_{1}\in T_{a_{1}},t_{2}\in T_{a_{2}},t\in T$, the following holds: